# Kähler compactification of $\mathbb{C}^{n}$ and Reeb dynamics

### v1

@@ 3.2. Symplectic homology and its variants
Let $W$ be the Liouville filling of $M$ in $\mathbb{C}^{n}$ in 2.8. To every Liouville filling, one can associate with several symplectic invariants, namely the symplectic homology (over $\mathbb{Q}$) $SH_{*}(W;\mathbb{Q})$ as well as its variants, see e.g. [CO18]. Those invariants are always graded over $\mathbb{Z}_{2}$, they are generated by (non-degenerate) Reeb orbits/Hamiltonian orbits with grading from the Conley-Zehnder index. When $c_{1}(W)=0$, those invariants are graded by $\mathbb{Z}$, where the grading depends on a choice of the trivialization of $\det_{\mathbb{C}}TW$. If we only have $c^{\mathbb{Q}}_{1}(W)=0$, those invariants can be graded over $\mathbb{Q}$ with similar dependence on trivializations.
@@
There are many different conventions regarding symplectic (co)homology and its variants. Since the references cited here is not completely consistent, we point out their relations: the symplectic homology $SH_{*}$ and its variants in [BO13, CO18, KvK16] is isomorphic to the symplectic cohomology $SH^{*}$ and its variants in [Rit13, Sei08] by $SH^{*}=SH_{n-*}$, where $n$ is half of the dimension of the symplectic domain. The symplectic homology in [BO13, CO18, KvK16] is graded by Conley-Zehnder indices of the Hamiltonian orbits.
@@
Let $W$ be the strongly pseudo-convex domain bounded by $M$ in $\mathbb{C}^{n}$ by 2.5, the positive $S^{1}$-equivariant symplectic cohomology of $W$ is
$SH^{+,S^{1}}_{*}(W;\mathbb{Q})=\left\{\begin{array}[]{cc}\mathbb{Q},&*=n+1+2m,m\in\mathbb{Z}_{\geq 0}\\ 0,&\text{else}\end{array}\right.$
@@
By gluing the Liouville filling $W$ from 2.8 and the Liouville cobordism in 2.9, we get a Liouville filling $\tilde{W}$ of the standard contact sphere. Then by a Theorem of Seidel and Smith [Sei08, Corollary 6.5], the symplectic homology of $\tilde{W}$ vanishes. We get a unital ring map $SH_{*}(\tilde{W};\mathbb{Q})\to SH_{*}(W;\mathbb{Q})$ from the Viterbo transfer [Rit13, Theorem 9.5], then we have $SH_{*}(W;\mathbb{Q})=0$. Then by the tautological long exact sequence [Rit13, Lemma 8.1]
@@
we get that $SH^{+}_{*}(W;\mathbb{Q})$ is supported in degree $n+1$ with rank $1$ as $W$ is a homology ball by 2.5. By [BO13, Theorem 1.1], we have the following Gysin exact sequence,
@@
As a consequence, we have $D:SH^{+,S^{1}}_{k}(W;\mathbb{Q})\to SH^{+,S^{1}}_{k-2}(W)$ is an isomorphism when $k\geq n+3$ or $k\leq n$. The map $D$ is nilpotent in the sense that for every $x\in SH^{+,S^{1}}_{*}(W;\mathbb{Q})$ there exists a $K\in\mathbb{Z}{\geq 0}$ depending on $x$, such that $D^{k}x=0$. This follows from the definition of $D$ [BO13] as it decreases the action. Therefore we have $SH^{+,S^{1}}_{k}(W;\mathbb{Q})=0$ when $k\leq n$. Then from the Gysin exact sequence, we have
@@
$\ldots\to 0\to SH^{+,S^{1}}_{n+2}(W;\mathbb{Q})\stackrel{{\scriptstyle D}}{{\to}}0\to\ldots,$
That is $SH^{+,S^{1}}_{n+1}(W;\mathbb{Q})=\mathbb{Q}$ and $SH^{+,S^{1}}_{n+2}(W;\mathbb{Q})=0$. Then the claim follows from $D$ is an isomorphism when $k\geq n+3$. ∎
The next result works for more general $M$, namely those 3.3 applies. Let $W$ be a Liouville symplectic filling of $M$ such that the rational first Chern class $c^{\mathbb{Q}}_{1}(W)=0$, we have the following Morse-Bott spectral sequence [KvK16, Theorem 5.4. (13)] computing $SH^{+,S^{1}}_{*}(W;\mathbb{Q})$.
$E_{p,q}^{1}=\bigoplus_{p=N(\ell+\frac{k}{|G|})}H_{p+q-\mu_{\mathrm{LCZ}}(G,k,\ell)}(Y_{G};\mathbb{Q})$
where $N=\mathrm{lcm}_{G\in S}(|G|)$. Here the spectral sequence is from the filtration by the period recorded by $p$ and $p+q$ is the (rational) grading on $SH^{+,S^{1}}_{*}(W;\mathbb{Q})$ by the (rational) Conley-Zehnder indices. This spectral sequence is induced by the increasing filtration by subcomplexes generated Reeb orbits with period up to $p/N$. This is a homological spectral sequence with differential $d^{r}_{p,q}:E^{r}_{p,q}\to E^{r}_{p-r,q+r-1}$. Moreover, each element from $H_{*}(\mathcal{Y}_{G};\mathbb{Q})$ also has a well-defined $\mathbb{Z}_{2}$ grading corresponding to the same spectral sequence from the period filtration but graded by the mod $2$ Conley-Zehnder index. The differential $d^{r}$ is degree $-1$ for both $\mathbb{Q}$-grading and $\mathbb{Z}_{2}$-grading. In the following picture, we spell out the $\mathbb{Q}$-grading and use different colors to mark the $\mathbb{Z}_{2}$-grading of each element.
$d^{1}$$d^{1}$$d^{2}$$p$$q$
@@
To present the first page of the spectral sequence, the vertical axis is $p$ and the horizontal axis is $q$, each dot represents a copy of $\mathbb{Q}$ with color indicating their $\mathbb{Z}_{2}$ grading. Each dashed block represents some $H_{*}(\mathcal{Y}_{G};\mathbb{Q})$ with the $p+q$ coordinate of the leading term is $\mu_{\mathrm{LCZ}}(G,k,\ell)$, all the leading dots have the same $\mathbb{Z}_{2}$ grading by (2d) of 3.3, where red means $n-1\mod 2$. The arrows are differentials that respect both gradings.
@@
Under the assumption of 3.3 and $R(M)>0$. If we have a Liouville filling $W$ of $Y$ such that $c_{1}^{\mathbb{Q}}(W)=0$, we have
$\mathrm{hmi}(M)=\inf\{k|SH^{+,S^{1}}_{k}(W;\mathbb{Q})\neq 0\}+n-3=\inf_{\gamma}(\mathrm{lSFT}(\gamma)),$
where $\gamma$ ranges over Reeb orbits of $\eta$.
@@
Since $SH^{+,S^{1}}_{*}(W;\mathbb{Q})$ is an invariant independent of $\eta$ whose underlying cochain complex is supported in degrees at least $\inf_{\gamma}(\mathrm{lSFT}(\gamma))+3-n$ for any contact form $\eta$. Therefore for any contact form $\eta$, we have
@@
where we ennumerate $\gamma$ through all Reeb orbits of $\eta$. Hence we know that
@@
Now we return to the conical contact form case. As $R(M)>0$, $\inf_{\gamma}(\mathrm{lSFT}(\gamma))=\min_{G,k}\{\mathrm{RS}(G,k,0)-\dim_{\mathbb{C}}Y_{G}+n-3\}\cup\{R(M)-2\}$. We claim that an element in $H_{0}(\mathcal{Y}_{G};\mathbb{Q})$ with the minimum rational degree $p+q$ and the maximum $p$ will survive in the spectral sequence. For such an element $\alpha$ to be killed in the spectral sequence, we must have an element $\beta$ in $E^{1}_{p+r,q-r+1}$ that survives till the $r$th page and $d^{r}\beta=\alpha$.
@@
Since the $\alpha$ and $\beta$ have different $\mathbb{Z}_{2}$ grading, we must have $\alpha$ red and $\beta$ blue, i.e. in the situation shown in the picture above. As a consequence, the minimal degree term $\gamma$ of the block containing $\beta$ must have strictly larger $p$ and no larger $p+q$ compared to $\alpha$, this contradicts with the choice of $\alpha$. Therefore $\alpha$ is non-trivial in $SH^{*}_{+,S^{1}}(W;\mathbb{Q})$ with the minimal grading $\inf_{\gamma}(\mathrm{lSFT}(\gamma))+3-n$. On the other hand, we have $\mathrm{hmi}(M)\geq\inf_{\gamma}(\mathrm{lSFT}(\gamma))$ by definition. Therefore we must achieve equality that
$\mathrm{hmi}(M)=\inf\{k|SH^{k}_{+,S^{1}}(W;\mathbb{Q})\neq 0\}+n-3=\inf_{\gamma}(\mathrm{lSFT}(\gamma))$
@@
Assume the Reeb flow of $\eta$ gives an $S^{1}$-action as above and $R(M)>0$. It is always true that $\mathrm{hmi}(M)$ is the $\inf_{\gamma}(\mathrm{lSFT}(\gamma))$, which is observed by a homology class, i.e. we do not need to assume the existence of a Liouville filling $W$. The existence of a Liouville filling allows us to use $\mathbb{Q}$-graded positive $S^{1}$-equivariant symplectic homology, which is the linearized contact homology w.r.t. the augmentation from $W$ [BO09]. In general, we can use the linearized contact homology w.r.t. an algebraic $\mathbb{Q}$-graded augmentation over $\mathbb{C}$. The Morse-Bott spectral sequence works for such linearized contact homology as well and the proof of 3.7 holds. As contact manifolds with $S^{1}$-Reeb flow are strongly fillable by [KL21], the contact homology does not vanish [LW11, Theorem 1 and Proposition 2.9]. Since the vanishing of contact homology is the only obstruction to the existence of an algebraic augmentation over an algebraically closed field, we can always use the linearized contact homology, which is expected to be a contact invariant [^3] The foundation of linearized contact homology is an ongoing project [ABZ] of Russell Avdek, Erkao Bao and the second author..
@@ 3.3. Proofs of the main theorems
It follows from (14), Propositions 3.4 and 3.7. ∎
Proposition 3.9.
The minimal discrepancy number of the cone singularity $\mathcal{C}(Y,\mathcal{L})$ is $n-1$.
Proof.
By Propositions 2.5 and 3.2, we have $R(M)>0$. Then Propositions 3.7 and 3.6 imply that $\inf_{\gamma}(\mathrm{lSFT}(\gamma))=\inf\{k|SH^{+,S^{1}}_{k}(W;\mathbb{Q})\neq 0\}+n-3=2n-2$, which implies the minimal discrepancy number is $n-1$ by 1.4. ∎
@@
When $\mathcal{Y}$ has no orbifold singularity, then Propositions 3.3 and 3.9 imply that $R(M)=n$. By 3.2, we have the Fano index $r(\mathcal{Y})=n$. Then the claim follows from [BM78, KO73]. ∎
Proof of Theorem 1.3.
By approximation, we can assume that there exists a sequence of rational ${\mathfrak{v}}_{k}$ that converges to ${\mathfrak{v}}$ as $k\rightarrow+\infty$. For each ${\mathfrak{v}}_{k}$, we apply Theorem 2.4 to get Kähler orbifold compactification of $\mathbb{C}^{n}$. By Proposition 2.10 we know that $\mathcal{C}$ has a Gorenstein isolated singularity at the vertex $o$.
We assume that Shokurov’s conjecture is true. Then $\mathcal{C}$ is a smooth affine variety with an effective torus action with a unique fixed attractive point $o\in\mathcal{C}$. By [KR82], we know that $\mathcal{C}$ is equivariant biholomorphic to $\mathbb{C}^{n}$ with a linear torus action. So the Reeb vector field of $(\mathcal{C},g_{0})\cong(\mathbb{C}^{n},g_{0})$ is exactly given by the imaginary part of some linear vector field $\sum_{i=1}^{n}w_{i}z_{i}\partial_{z_{i}}$ with $w_{i}>0,i=1,\dots,n$. Now if $g$ is Ricci flat, then $(\mathcal{C},g_{0})$ is also Ricci flat. By the volume minimization property of Reeb vector fields associated to Ricci flat Kähler cone metrics (see [MSY08]), the vector $(w_{1},\dots,w_{n})$ must minimize the normalized functional $\frac{(w_{1}+\cdots+w_{n})^{n}}{w_{1}\cdot\cdots\cdot w_{n}}$. This forces $w_{1}=w_{2}=\cdots=w_{n}$ and $g_{0}$ is the flat metric on $\mathbb{C}^{n}$. In other words, the metric tangent cone of $(\mathbb{C}^{n},g)$ at infinity is the flat metric on $\mathbb{C}^{n}$. By the rigidity result from [And90], we know that the Ricci flat metric $g$ on $\mathbb{C}^{n}$ itself must be flat. ∎
@@
$\pi^{-1}(U)=\mathrm{Spec}(\mathbb{C}[x_{1}^{m},x_{2}x_{1}^{-e_{2}},\dots,x_{n}x_{1}^{-e_{n}}]).$
Assume that $-K_{\mathcal{Y}}=r\mathcal{L}$ so that $\mathcal{C}$ is $\mathbb{Q}$-Gorenstein, i.e. there exists $N\in\mathbb{Z}_{>0}$ such that $N\cdot K_{\mathcal{C}}$ is Cartier. Then we can write down a local trivializing section of $(K_{\mathcal{C}})^{\otimes N}$ as ${\sigma}_{\mathcal{C}}=\left(x_{1}^{r-1}dx_{1}\wedge dx_{2}\wedge\cdots\wedge dx_{n}\right)^{\otimes N}$ which induces a trivializing section of $((\det\xi)^{\vee})^{\otimes N}$ as ${\sigma}:=(\iota_{x_{1}\partial_{x_{1}}}\sigma_{\mathcal{C}})^{\otimes N}=(x_{1}^{r}dx_{2}\wedge\cdots\wedge dx_{n})^{\otimes N}$. Choose an orbifold Hermitian metric $h$ on the orbifold line bundle $\mathcal{L}^{-1}\rightarrow\mathcal{Y}$ such that its Chern curvature $\sqrt{-1}\partial\bar{\partial}\log h$ is an orbifold Kähler form. Locally over the uniformization chart $h$ is represented as $a(x^{\prime})|x_{1}|^{2}$ where $x^{\prime}=\{x_{2},\dots,x_{n}\}$. For any $\beta>0$, $h^{\beta}$ induces a radius function $\mathfrak{r}$ on $\mathcal{C}$ which in turn defines a Riemannian cone metric $g_{\mathcal{C}}$ on $\mathcal{C}$. Over the regular part $\mathcal{C}^{\circ}$, we can find a basis of the contact distribution $\xi$ that are compatible with the trivialization of $(\det\xi)^{\otimes N}$:
@@
Indeed, it is easy to verify that $\{V_{j}:2\leq j\leq n\}$ are basis for the orthogonal complement to the Euler vector field $x_{1}\partial_{x_{1}}$ and are invariant under the cyclic group action. Let $\nabla$ denote the Levi-Civita connection of $g_{\mathcal{C}}|_{\{\mathfrak{r}=1\}}$. Set $x_{1}=|x_{1}|e^{\sqrt{-1}\theta}$. By a straightforward calculation, we get:
@@
So if we choose $\beta=\frac{r}{n-1}$ then the parallel transport trivializes $(\det\xi)^{\otimes N}$. From now on, we set $\beta$ to be of this value. Let $\gamma$ be the parametrised closed Reeb orbit: $\theta\rightarrow[(\theta,0,\dots,0)]$ with $\theta\in[0,\frac{2\pi}{m}]$. We now write down a trivialization of $\gamma^{*}\xi^{\oplus N}$ by an appropriate choice of basis that is compatible with the trivialization of $\det(\xi)^{\otimes N}$. First, we consider the Gorenstein case i.e. when $N=1$. We set

### v3

@@ 3.2. Symplectic homology and its variants
Let $W$ be the Liouville filling of $M$ in $\mathbb{C}^{n}$ in 2.13. To every Liouville filling, one can associate several symplectic invariants, namely the symplectic homology (over $\mathbb{Q}$) $SH_{*}(W;\mathbb{Q})$ as well as its variants, see e.g. [CO18]. Those invariants are always graded over $\mathbb{Z}_{2}$, and they are generated by (non-degenerate) Reeb orbits/Hamiltonian orbits with grading from the Conley-Zehnder index. When $c_{1}(W)=0$, those invariants are graded by $\mathbb{Z}$, where the grading depends on a choice of the trivialization of $\det_{\mathbb{C}}TW$. If we only have $c^{\mathbb{Q}}_{1}(W)=0$, those invariants can be graded over $\mathbb{Q}$ with similar dependence on trivializations.
@@
There are many different conventions regarding symplectic (co)homology and its variants. Since the references cited here are not completely consistent, we point out their relations: the symplectic homology $SH_{*}$ and its variants in [BO13, CO18, KvK16] is isomorphic to the symplectic cohomology $SH^{*}$ and its variants in [Rit13, Sei08] by $SH^{*}=SH_{n-*}$, where $n$ is half of the dimension of the symplectic domain. The symplectic homology in [BO13, CO18, KvK16] is graded by Conley-Zehnder indices of the Hamiltonian orbits.
@@
Let $W$ be the strongly pseudo-convex domain bounded by $M$ in $\mathbb{C}^{n}$ obtained by 2.10, the positive $S^{1}$-equivariant symplectic homology of $W$ is
$SH^{+,S^{1}}_{*}(W;\mathbb{Q})=\left\{\begin{array}[]{cc}\mathbb{Q},&*=n+1+2m,m\in\mathbb{Z}_{\geq 0}\\ 0,&\text{else}\end{array}\right.$.
@@
By gluing the Liouville filling $W$ from 2.13 and the Liouville cobordism in 2.14, we get a Liouville filling $\tilde{W}$ of the standard contact sphere. Then by a theorem of Seidel and Smith [Sei08, Corollary 6.5], the symplectic homology of $\tilde{W}$ vanishes. Because there is a unital ring map $SH_{*}(\tilde{W};\mathbb{Q})\to SH_{*}(W;\mathbb{Q})$ from the Viterbo transfer [Rit13, Theorem 9.5], we have the vanishing $SH_{*}(W;\mathbb{Q})=0$. Then by the tautological long exact sequence [Rit13, Lemma 8.1]
@@
we get that $SH^{+}_{*}(W;\mathbb{Q})$ is supported in degree $n+1$ with rank $1$ as $W$ is a homology ball by 2.10. By [BO13, Theorem 1.1], we have the following Gysin exact sequence,
@@
As a consequence, we know that $D:SH^{+,S^{1}}_{k}(W;\mathbb{Q})\to SH^{+,S^{1}}_{k-2}(W)$ is an isomorphism when $k\geq n+3$ or $k\leq n$. The map $D$ is nilpotent in the sense that for every $x\in SH^{+,S^{1}}_{*}(W;\mathbb{Q})$ there exists a $K\in\mathbb{Z}{\geq 0}$ depending on $x$, such that $D^{k}x=0$. This follows from the definition of $D$ [BO13] as it decreases the action. Therefore we have $SH^{+,S^{1}}_{k}(W;\mathbb{Q})=0$ when $k\leq n$. Then from the Gysin exact sequence, we have exact sequences:
@@
$\ldots\to 0\to SH^{+,S^{1}}_{n+2}(W;\mathbb{Q})\stackrel{{\scriptstyle D}}{{\to}}0\to\ldots.$
So we get $SH^{+,S^{1}}_{n+1}(W;\mathbb{Q})=\mathbb{Q}$ and $SH^{+,S^{1}}_{n+2}(W;\mathbb{Q})=0$. The claim now follows because we know that $D$ is an isomorphism when $k\geq n+3$. ∎
The next result works for more general $M$, namely for those 3.3 applies. Let $W$ be a Liouville symplectic filling of $M$ such that the rational first Chern class $c^{\mathbb{Q}}_{1}(W)=0$. We have a Morse-Bott spectral sequence [KvK16, Theorem 5.4. (5.3)] computing $SH^{+,S^{1}}_{*}(W;\mathbb{Q})$. Its first page is given by:
$E_{p,q}^{1}=\bigoplus_{p=N(\ell+\frac{k}{|G|})}H_{p+q-\mu_{\mathrm{LCZ}}(G,k,\ell,i)}(\mathcal{Y}^{i}_{G};\mathbb{Q})$
where $N=\mathrm{lcm}_{G\in S}(|G|)$. Here, the spectral sequence is from the filtration by the period recorded by $p$ and $p+q$ is the (rational) grading on $SH^{+,S^{1}}_{*}(W;\mathbb{Q})$ by the (rational) Conley-Zehnder indices. This spectral sequence is induced by the increasing filtration by subcomplexes generated by Reeb orbits with period up to $p/N$. This is a homological spectral sequence with differential $d^{r}_{p,q}:E^{r}_{p,q}\to E^{r}_{p-r,q+r-1}$. Moreover, each element from $H_{*}(\mathcal{Y}^{i}_{G};\mathbb{Q})$ also has a well-defined $\mathbb{Z}_{2}$ grading corresponding to the same spectral sequence from the period filtration but graded by the mod $2$ Conley-Zehnder index. The differential $d^{r}$ is degree 1-1 for both the $\mathbb{Q}$-grading and $\mathbb{Z}_{2}$-grading. In the following picture, we spell out the $\mathbb{Q}$-grading and use different colors to mark the $\mathbb{Z}_{2}$-grading of each element.
Remark 3.7.
[KvK16, Theorem 5.4, condition (ii)] requires that $c_{1}(\xi),\pi_{1}(M)=0$. Such a condition is only needed to obtain a canonical $\mathbb{Z}$-grading on symplectic homology. Moreover, the actual condition needed to obtain a canonical $\mathbb{Z}$-grading is that all Reeb orbits are null-homologous and $c_{1}(\xi)=0$. While the spectral sequence is induced from a period/action filtration, which always exists in a Morse-Bott situation by the proof of [KvK16, Theorem 5.4]. The extra condition equips the spectral sequence with $\mathbb{Z}$-grading. In our case here, we have $c_{1}^{\mathbb{Q}}(\xi)=0$ and all Reeb orbits have torsion homology class, this equips symplectic homology and the spectral sequence with a canonical $\mathbb{Q}$-grading using the rational Conley-Zehnder index recalled in §2.3.
$d^{1}$$d^{1}$$d^{2}$ppqq
@@
To present the first page of the spectral sequence, the vertical axis is $p$ and the horizontal axis is $q$; each dot represents a copy of $\mathbb{Q}$ with color indicating their $\mathbb{Z}_{2}$ grading. Each dashed block represents some $H_{*}(\mathcal{Y}^{i}_{G};\mathbb{Q})$ with the $p+q$ coordinate of the leading term is $\mu_{\mathrm{LCZ}}(G,k,\ell,i)$. All the leading dots have the same $\mathbb{Z}_{2}$ grading by (2d) of 3.3, where red means $n-1\mod 2$. The arrows are differentials that respect both gradings.
@@
Under the assumption of 3.3 and $R(M)>0$. If there is a Liouville filling $W$ of $M$ such that $c_{1}^{\mathbb{Q}}(W)=0$, we have
$\inf_{\gamma}(\mathrm{lSFT}(\gamma))=\inf\{k|SH^{+,S^{1}}_{k}(W;\mathbb{Q})\neq 0\}+n-3=\mathrm{hmi}(M),$
where $\gamma$ ranges over Reeb orbits of a conic contact form $\eta$.
@@
We know that $SH^{+,S^{1}}_{*}(W;\mathbb{Q})$ is an invariant independent of $\eta$ whose underlying cochain complex is supported in degrees at least $\inf_{\gamma}(\mathrm{lSFT}(\gamma))+3-n$ for any contact form $\eta$. Therefore for any contact form $\eta$, we have
@@
where we enumerate $\gamma$ through all Reeb orbits of $\eta$. Hence, we know that
@@
Now we return to the case of conic contact forms. As $R(M)>0$, $\inf_{\gamma}(\mathrm{lSFT}(\gamma))=\min_{G,k}\{\mu_{\mathrm{RS}}(G,k,0,i)-\dim_{\mathbb{C}}\mathcal{Y}^{i}_{G}+n-3\}\cup\{2R(M)-2\}$. We claim that an element in $H_{0}(\mathcal{Y}^{i}_{G};\mathbb{Q})$ with the minimal rational degree $p+q$ and the maximal $p$ will survive in the spectral sequence. The $R(M)>0$ condition implies that concatenation with principal orbits increases both the period and grading, hence such elements exist. For such an element $\alpha$ to be killed in the spectral sequence, we must have an element $\beta$ in $E^{1}_{p+r,q-r+1}$ that survives till the $r$th page and $d^{r}\beta=\alpha$.
@@
Since the $\alpha$ and $\beta$ have different $\mathbb{Z}_{2}$ grading, we must have $\alpha$ red and $\beta$ blue, i.e. in the situation shown in the picture above. As a consequence, the minimal degree term $\gamma$ of the block containing $\beta$ must have strictly larger $p$ and no larger $p+q$ compared to $\alpha$. But this contradicts with the choice of $\alpha$. Therefore $\alpha$ is non-trivial in $SH_{*}^{+,S^{1}}(W;\mathbb{Q})$ with the minimal grading $\inf_{\gamma}(\mathrm{lSFT}(\gamma))+3-n$. On the other hand, we have $\mathrm{hmi}(M)\geq\inf_{\gamma}(\mathrm{lSFT}(\gamma))$ by definition. So we must achieve the following equalities:
$\inf_{\gamma}(\mathrm{lSFT}(\gamma))=\inf\{k|SH_{k}^{+,S^{1}}(W;\mathbb{Q})\neq 0\}+n-3=\mathrm{hmi}(M).$
@@
Assume the Reeb flow of $\eta$ gives an $S^{1}$-action as above and $R(M)>0$. It is always true that $\mathrm{hmi}(M)$ is the $\inf_{\gamma}(\mathrm{lSFT}(\gamma))$, which is observed by a homology class, i.e. we do not need to assume the existence of a Liouville filling $W$. The existence of a Liouville filling allows us to use $\mathbb{Q}$-graded positive $S^{1}$-equivariant symplectic homology, which is the linearized contact homology w.r.t. the augmentation from $W$ [BO09]. Such a Liouville filling always exists in our setting if $o\in\mathcal{C}$ is smoothable and $\dim_{\mathbb{C}}\mathcal{C}\geq 3$, or in the compactification problem by 2.13. In general, we can use the linearized contact homology w.r.t. an algebraic $\mathbb{Q}$-graded augmentation over $\mathbb{C}$. The Morse-Bott spectral sequence works for such linearized contact homology as well, and the proof of 3.8 holds. As contact manifolds with $S^{1}$-Reeb flow are strongly fillable by [KL21], the contact homology does not vanish [LW11, Theorem 1 and Proposition 2.9]. Since the vanishing of contact homology is the only obstruction to the existence of an algebraic augmentation over an algebraically closed field, we can always use the linearized contact homology, which is expected to be a contact invariant [^5] The foundation of linearized contact homology is an ongoing project [ABDRZ] of Russell Avdek, Erkao Bao, Georgios Dimitroglou Rizell, and the second author..
@@ 3.3. Proofs of the main theorems
The statement follows immediately by combining (15), Propositions 3.4 and Proposition 3.8. ∎
Proof of Theorem 1.3.
By approximation, we can assume that there exists a sequence of quasi-regular ${\mathfrak{v}}_{k}$ that converges to $\mathfrak{v}$} as $k\rightarrow+\infty$. For each ${\mathfrak{v}}_{k}$, we apply Theorem 2.7 to get Kähler orbifold compactification of $\mathbb{C}^{n}$. By Proposition 2.15 we know that $\mathcal{C}$ has a Gorenstein isolated singularity at the vertex $o$.
By Propositions 2.10 and Proposition 3.2, we have $R(M)>0$. Then Propositions 3.8 and Proposition 3.6 imply that $\inf_{\gamma}(\mathrm{lSFT}(\gamma))=\inf\{k|SH^{+,S^{1}}_{k}(W;\mathbb{Q})\neq 0\}+n-3=2n-2$, which implies the minimal discrepancy number is n1n-1 by 1.4.
If we assume that Shokurov’s conjecture is true, in particular if the dimension $n\leq 3$, then $\mathcal{C}$ is a smooth affine variety with an effective torus action with a unique fixed attractive point $o\in\mathcal{C}$. By [KR82], we know that $\mathcal{C}$ is equivariant biholomorphic to $\mathbb{C}^{n}$ with a linear torus action. So the Reeb vector field of $(\mathcal{C},g_{0})\cong(\mathbb{C}^{n},g_{0})$ is exactly given by the imaginary part of some linear vector field $\sum_{i=1}^{n}w_{i}z_{i}\partial_{z_{i}}$ with $w_{i}>0,i=1,\dots,n$. Now if $g$ is Ricci flat, then $(\mathcal{C},g_{0})$ is also Ricci flat. By the volume minimization property of Reeb vector fields associated to Ricci flat Kähler cone metrics (see [MSY08]), the vector $(w_{1},\dots,w_{n})$ must minimize the normalized functional $\frac{(w_{1}+\cdots+w_{n})^{n}}{w_{1}\cdot\cdots\cdot w_{n}}$. This forces $w_{1}=w_{2}=\cdots=w_{n}$ and $g_{0}$ is the flat metric on $\mathbb{C}^{n}$. In other words, the metric tangent cone of $(\mathbb{C}^{n},g)$ at infinity is the flat metric on $\mathbb{C}^{n}$. By the rigidity result from [And90], we know that the Ricci flat metric $g$ on $\mathbb{C}^{n}$ itself must be flat. ∎
@@
By [BM78, Theorem 1.1.III] or Proposition 2.10, we know that $X$ and $Y$ are Fano manifolds satisfying $K_{X}=(r+1)[Y]$ for $r>0$. According to the construction of Tian-Yau (see Example 2.6), there exists a complete AC metric on $X\setminus Y$ whose asymptotical cone is $\mathcal{C}(Y,L)$ with $L=[Y]|_{Y}$. So we can apply Theorem 1.3 to the case when $(\mathcal{Y},\mathcal{L})=(Y,L)$ has no orbifold singularities. Since $K_{Y}=rL$, by (15) it is easy to see that $\mathrm{md}(o,\mathcal{C}(Y,L))+1$ is now equal to $r$. So by Theorem 1.3, $r$ is equal to $n$. By Kobayashi-Ochiai’s criterion ([KO73]), we know that $Y$ is holomorphic to $\mathbb{P}^{n-1}$ and the conclusion now follows from the results in [BM78, Theorem 1.1.(II)]. ∎
@@
$\mathbb{C}^{*}\times\mathbb{C}^{n-1}/\frac{1}{m}(1,e_{2},\dots,e_{n})=\mathrm{Spec}(\mathbb{C}[x_{1}^{m},x_{2}x_{1}^{-e_{2}},\dots,x_{n}x_{1}^{-e_{n}}]).$
Assume that $K_{\mathcal{Y}}=r\mathcal{L}$ so that $\mathcal{C}$ is $\mathbb{Q}$-Gorenstein, i.e. there exists $N\in\mathbb{Z}_{>0}$ such that $N\cdot K_{\mathcal{C}}$ is Cartier. Then we can write down a local trivializing section of $(K_{\mathcal{C}})^{\otimes N}$ as
${\sigma}_{\mathcal{C}}=\left(x_{1}^{r-1}dx_{1}\wedge dx_{2}\wedge\cdots\wedge dx_{n}\right)^{\otimes N}$
which induces a trivializing section of $((\det\xi)^{\vee})^{\otimes N}$ as ${\sigma}:=(\iota_{x_{1}\partial_{x_{1}}}\sigma_{\mathcal{C}})^{\otimes N}=(x_{1}^{r}dx_{2}\wedge\cdots\wedge dx_{n})^{\otimes N}$. Choose an orbifold Hermitian metric $h$ on the orbifold line bundle $\mathcal{L}^{-1}\rightarrow\mathcal{Y}$ such that its Chern curvature $\sqrt{-1}\partial\bar{\partial}\log h$ is an orbifold Kähler form. Locally over the uniformization chart $h$ is represented as $a(x^{\prime})|x_{1}|^{2}$ where $x^{\prime}=\{x_{2},\dots,x_{n}\}$. For any $\beta>0$, $h^{\beta}$ induces a radius function $\mathfrak{r}$ on $\mathcal{C}$ which in turn defines a Riemannian cone metric $g_{\mathcal{C}}$ on $\mathcal{C}$. Over the regular part $\mathcal{C}^{\circ}$, we can find a basis of the contact distribution $\xi$ that are compatible with the trivialization of $(\det\xi)^{\otimes N}$:
@@
Indeed, it is easy to verify that $\{V_{j}:2\leq j\leq n\}$ are a basis for the orthogonal complement to the Euler vector field $x_{1}\partial_{x_{1}}$ and are invariant under the cyclic group action. Let $\nabla$ denote the Levi-Civita connection of $g_{\mathcal{C}}|_{\{\mathfrak{r}=1\}}$. Set $x_{1}=|x_{1}|e^{\sqrt{-1}\theta}$. By a straightforward calculation, we get:
@@
So if we choose $\beta=\frac{r}{n-1}$ then the parallel transport trivializes $(\det\xi)^{\otimes N}$. From now on, we set $\beta$ to be of this value. Let $\gamma$ be the parametrized closed Reeb orbit: $\theta\rightarrow[(\theta,0,\dots,0)]$ with $\theta\in[0,\frac{2\pi}{m}]$. We now write down a trivialization of $\gamma^{*}\xi^{\oplus N}$ by an appropriate choice of basis that is compatible with the trivialization of $\det(\xi)^{\otimes N}$. First, we consider the Gorenstein case, i.e. when $N=1$. We set